OpenStudy (anonymous):

f(x)=x^2 +1 find (goh) that is equal to f(x)

OpenStudy (anonymous):

www.geogebra.org math troubles = over. :)

OpenStudy (anonymous):

wot?

OpenStudy (anonymous):

it asks for g(x) and h(x) that (goh) is equal to f(x)

OpenStudy (anonymous):

h(t) = t^2 g(x) = x+1

OpenStudy (anonymous):

Haha. :) No, I meant if you go to that site, your math troubles will be over. xD Here's the direct link: http://www.geogebra.org/cms/

OpenStudy (anonymous):

thank you guys...tell me about the software

OpenStudy (anonymous):

It helps you solve simple math equations and you don't have to download either. :)

OpenStudy (anonymous):

is it free?

OpenStudy (anonymous):

Just click "download" and then "applet start". It will open up in a seperate web browser. Yes, 100% free.

OpenStudy (anonymous):

webstart or geogebraprim?

OpenStudy (anonymous):

"webstart" if you want to download. :)